OpenStudy (anonymous):

Line AB contains points A (0, 0) and B (2, 2). Line CD contains points C (3, 1) and D (5, 3). Lines AB and CD are ???

OpenStudy (kira_yamato):

First we find gradient of the lines using the formula \[m = \frac{y_1 - y_2}{x_1 - x_2}\] m(AB) = 1 m(CD) = 1 Since m(AB) = m(CD) the lines are parallel